# Review: Quantum Metrology and Sensing with Many-Body Systems

### v1

@@ I Introduction
The rapidly emerging quantum technologies are expected to revolutionize our lives in the coming decades Dowling and Milburn (2003). In a broad sense, quantum technologies are developed in three directions, namely quantum simulation and computation Lloyd (1996); Georgescu et al. (2014), quantum communication Zeilinger (1999); Gisin and Thew (2007), and quantum sensing Degen et al. (2017). Thanks to advancements in fabricating ultra-precise quantum devices, the progress in all these three aspects have significantly been accelerated. Quantum sensing Degen et al. (2017) is one of the pillars of quantum technologies which allow for developing a new generation of sensors for detecting gravitational, magnetic and electric fields as well as estimating physical quantities with unprecedented precision, well beyond the capability of conventional classical sensors. The application prospect of quantum sensors is very wide and includes mining Crawford et al. (2021), environmental monitoring  Lanzagorta et al. (2015), biological imaging Wu et al. (2016a); Taylor and Bowen (2016); Aslam et al. (2023), space exploration Kaltenbaek et al. (2021), radar technology Maccone and Ren (2020); Karsa et al. (2024), searching for fundamental particles Brady et al. (2022); Backes et al. (2021); Kang et al. (2024); Wang et al. (2022a); Jiang et al. (2021a) — the list goes on. There are several reasons that make quantum sensors superior over their classical counterparts. First, the even smaller sizes of quantum sensors, which can be fabricated at the atomic scales, allows for excellent spatial resolution Schirhagl et al. (2014); Zhang et al. (2021). Second, the natural delicacy of quantum superposition allows quantum sensors to detect very weak signals Hong et al. (2024); Jiang et al. (2021b). Third, a combination of quantum particles can benefit from a unique property of quantum physics, namely entanglement, which allows for quantum enhanced precision in comparison with a classical sensor using the same amount of resources Giovannetti et al. (2004, 2011, 2006).
@@
Quantum criticality has already been identified as a resource for sensing in both equilibrium and non-equilibrium many-body systems. Originally, quantum phase transition in the transverse Ising model has been explored for metrology purposes Zanardi and Paunković (2006). It has been then extended to various types of criticalities including, first-order Raghunandan et al. (2018); Heugel et al. (2019); Yang and Jacob (2019), second-order Zanardi and Paunković (2006); Zanardi et al. (2007); Gu et al. (2008); Zanardi et al. (2008); Invernizzi et al. (2008); Gu (2010); Gammelmark and Mølmer (2011); Skotiniotis et al. (2015); Rams et al. (2018); Wei (2019); Chu et al. (2021); Liu et al. (2021); Montenegro et al. (2021); Mirkhalaf et al. (2021); Di Candia et al. (2023), topological Budich and Bergholtz (2020); Sarkar et al. (2022); Koch and Budich (2022); Yu et al. (2024a); Sarkar et al. (2024), non-Hermitian systems Chen et al. (2017); Yu et al. (2020); McDonald and Clerk (2020); Wiersig (2014); Hodaei et al. (2017); Liu et al. (2016); Langbein (2018); Chen et al. (2019); Wang et al. (2020); Zhang et al. (2019); Lau and Clerk (2018); Kononchuk et al. (2022); Wiersig (2020); Xiao et al. (2024); Yu et al. (2024b), and Stark phase transitions He et al. (2023); Yousefjani et al. (2023a, 2024a). Apart from these distinct types of criticalities at equilibrium, non-equilibrium quantum phase transitions can also demonstrate quantum enhanced sensitivity. This includes dissipative Fernández-Lorenzo and Porras (2017); Baumann et al. (2010); Baden et al. (2014); Klinder et al. (2015); Rodriguez et al. (2017); Fitzpatrick et al. (2017); Fink et al. (2017); Ilias et al. (2022, 2024); Alipour et al. (2014), Floquet Mishra and Bayat (2021, 2022) and time crystal Montenegro et al. (2023); Cabot et al. (2024); Lyu et al. (2020); Iemini et al. (2024a); Yousefjani et al. (2024b) phase transitions. Each type of quantum criticality comes with its own features and characteristics. For instance, second-order quantum phase transitions take place in the ground state of a many-body system and are accompanied by spontaneous symmetry breaking and long-range correlations which are described by Landau-Ginzburg formalism Sachdev (1999). On the other hand, topological phase transitions neither show symmetry breaking nor follow the Landau-Ginzburg theorem Xu (2012) and Stark localization takes place across the entire spectrum and not just the ground state van Nieuwenburg et al. (2019); Schulz et al. (2019). While different types of quantum criticalities have their own characteristics, one feature is common among all of them for showing quantum enhanced sensitivity which is energy/quasi-energy gap closing. This is an interesting observation which guides us in searching for finding new potential quantum sensors with the capability of achieving quantum enhanced precision. Interestingly, the realization of criticality-based quantum sensors are becoming viable in various physical platforms, including nuclear magnetic resonance systems Liu et al. (2021) and Rydberg atoms Ding et al. (2022).
@@
Quantum criticality is in general a property of equilibrium states such as the Gibbs thermal state (which becomes the ground state at zero temperature limit) Sachdev and Safari (2011); Vojta (2003). Preparing such equilibrium states can be challenging in practice, demanding ultra-low temperatures or extremely long preparation times. In addition, criticality-enhanced quantum precision is only achievable around the phase transition point. This requires fine tuning and thus is mostly beneficial in the context of local sensing, where significant prior information about the parameter of interest is available. Therefore, as an alternative, one can use non-equilibrium dynamics of many-body systems for sensing unknown parameters. These setups are more experimental friendly as their initial state can be a simple product state and entanglement is generated naturally during the dynamics. Non-equilibrium dynamics in many-body systems can be induced through different methods, including quantum quench Mitra (2018); Daley et al. (2012); Calabrese and Cardy (2006); Calabrese et al. (2011), measurement quench Pouyandeh et al. (2014); Bayat et al. (2018); Skinner et al. (2019a); Li et al. (2018a) and external driving Chu and Telnov (2004); Gritsev and Polkovnikov (2017); Oka and Kitamura (2019). Apart from the probe size, in non-equilibrium quantum sensors, time is also a key resource which plays an important role and has to be included in the definition of standard limit and quantum enhanced sensitivity. In general, Fisher information scales as $I\sim t^{\alpha}L^{\beta}$, where $t$ is time, $L$ is the system size and $\alpha$ and $\beta$ are two exponents Ilias et al. (2022). In non-equilibrium sensors, the standard limit is defined as $\alpha{=}\beta{=}1$. Any situation with $\alpha>1$ or $\beta>1$ is called quantum enhanced sensitivity with the special case of Heisenberg limit, defined as $\alpha{=}\beta{=}2$. In certain systems in which the unitary dynamics preserves the number of excitations, the Fisher information may scale with the number of excitations too. In fact, in Ref. Manshouri et al. (2024) it was shown that in excitation-preserving dynamics Fisher information scales as $I\sim t^{\alpha}L^{\beta}N^{\gamma}$, where $N$ is the number of excitations and $\gamma$ is its corresponding exponent. This is an interesting observation which shows how the sensitivity increases from single excitation subspace, i.e. $N{=}1$, to half filling, i.e. $N{=}L/2$. Unlike criticality-based quantum sensors, the criteria for achieving quantum enhanced precision is not yet well-characterized in non-equilibrium quantum many-body probes. For instance, we know that localized phases are not good for sensing He et al. (2023); Yousefjani et al. (2023a, 2024a); Sahoo et al. (2024) but what features in extended phases are crucial for quantum enhanced sensitivity are not yet identified.
@@
In practice, no quantum device is perfectly isolated from its environment and thus decoherence is an inevitable part of any quantum protocol. Consequently, any quantum dynamics is actually an open system evolution. Different formalisms have been developed for addressing open system dynamics, among them are: (i) master equation Carmichael (1993); and (ii) non-Hermitian Hamiltonians Rotter (2009). In the former, the evolution of the system and its environment is described by a unitary operation. By tracing out the environment, the evolution of the system, typically within Born-Markov approximations, is thus characterized by the Lindblad (or Gorini-Kossakowski-Sudarshan-Lindblad) master equation Breuer et al. (2002). In this formalism a quantum system initially prepared in a pure state may indeed evolve to a mixed state due to the action of jump operators — a signature of information leakage (irreversible process) to the environment. In non-Hermitian formalism, the environment is monitored for tracking a trajectory in which no jump operator action takes place Plenio and Knight (1998); Daley (2014). As a result, an initially pure quantum state remains pure throughout the dynamics, despite being evolved effectively as an open quantum system. Other mechanisms that address or profit from dissipation include decoherence-free subspaces Lidar et al. (1998); Duan and Guo (1997), dynamical decoupling Viola et al. (1999); Khodjasteh and Lidar (2005), and quantum reservoir engineering Poyatos et al. (1996). The latter technique involves adding an engineered Lindbladian term to the open system dynamics. This approach can be designed to steer the system into a desired pure steady-state, effectively cooling the system to that state Myatt et al. (2000); Pielawa et al. (2007, 2010); Groszkowski et al. (2022). Both master equation and non-Hermitian formalisms have been investigated for quantum sensing. Indeed, a general joint (system plus environment) quantum Fisher information upper bound has been formulated Escher et al. (2011a). Hence, one would expect that such open quantum system dynamics may also achieve quantum enhanced sensitivity provided that the system goes through a phase transition, which is again accompanied by an eigenvalue gap closing. In the master equation formalism, the phase transition is determined for the steady state of the system at the point that the eigenvalues of the Liouvillian operator shows gap closing Minganti et al. (2018). An example of such systems with quantum enhanced sensitivity is boundary time crystals Montenegro et al. (2023). In the non-Hermitian systems, quantum enhanced sensitivity has been shown to emerge at the exceptional points Wiersig (2014); Kononchuk et al. (2022), where two or more eigenvalues and corresponding eigenstates coalesce, as well as in the cases of different types of gap closing in the complex energy plane Sarkar et al. (2024).
@@ II.1 Global and local estimation theory
A suitable strategy to estimate a parameter usually depends on the amount of a priori information one has on the parameter itself. If no a priori information is avalaible, the estimation strategy, i.e. the choice of the measure to be performed on the probe and the procedure employed to process data, should be optimal in average i.e. should be equally good for any value of the parameter. The set of techniques and methods used to choose and optimize the figure of merit in this case is known as global estimation theory Helstrom (1976); Holevo (2011a). Once a global estimation has been performed an experimenter is left with some a priori information to be improved, i.e. one should look for figures of merits that select the best estimation strategy for a specific value or range of values of a parameter. This optimization problem is addressed by local estimation theory Bengtsson and Życzkowski (2017); Van Trees (2004); Helstrom (1976) which allows one to achieve the ultimate bound on precision. The quantum version of local quantum estimation theory Braunstein and Caves (1994); Paris (2009) is the fundamental tool of quantum probing, sensing and metrology and is the main subject of the present review, with the exception of Section XIV, where various aspects of global sensing are discussed.
@@ II.3 Optimizing the classical bound: the QFI
Note that the derivative of the quantum state satisfies other expressions, e.g., the Right Logarithmic Derivative (RLD) Albarelli et al. (2020). This gives rise to different quantum lower bounds. Hence, as opposed to the classical Fisher information, the QFI is not unique (see Ref. Meyer (2021) for a thorough discussion and Ref. Albarelli et al. (2020) for different achievable bounds).
The QFI of Eq. (6) accounts for the ultimate precision limit for the estimation of $\theta$. Note that the QFI, as the maximization of the classical Fisher information over all possible measurements, depends on the quantum statistical model $\rho_{\theta}$ and the SLD $L_{\theta}$ only, and it does not depend explicitly on a particular measurement basis. The optimal POVM can be composed with the set of projectors over the eigenstates of $L_{\theta}$ with an associated optimal quantum estimatorParis (2009).
To saturate the quantum Cramér-Rao bound in Eq. (2), one needs both the optimal POVM (as determined by the support of $L_{\theta}$) and an optimal classical estimator for post-processing data, such as the maximum-likelihood estimator in the limit of large data sets Olivares and Paris (2009); Pezzé et al. (2007); Hradil et al. (1996); Le Cam (1986); Cramer (1946); Braunstein and Caves (1994); Helstrom (1967) (refer to Refs. Rubio and Dunningham (2019, 2020) for Bayesian analysis in the scenario of limited data). The optimal POVM ensures that the classical Fisher information equals the QFI, i.e., $I(\theta)=I_{Q}(\theta)$. This is crucial as it determines that the single-parameter quantum Cramér-Rao bound can always be achieved, provided that both the estimator and the measurement basis are optimal.
@@
The QFI of Eqs. (6)-(11) applies to any quantum state. However, for pure quantum states $\rho_{\theta}=|\psi_{\theta}\rangle\langle\psi_{\theta}|$ such that $\text{Tr}[\rho_{\theta}^{2}]=1$, the QFI adopts a simpler form given by Paris (2009); Sidhu and Kok (2020); Liu et al. (2019):

### v2

@@ I Introduction
The rapidly emerging quantum technologies are expected to revolutionize our lives in the coming decades Dowling and Milburn (2003). In a broad sense, quantum technologies are developed in three directions, namely quantum simulation and computation Lloyd (1996); Georgescu et al. (2014), quantum communication Zeilinger (1999); Gisin and Thew (2007), and quantum sensing Degen et al. (2017). Thanks to advancements in fabricating ultra-precise quantum devices, the progress in all these three aspects have significantly been accelerated. Quantum sensing Degen et al. (2017) is one of the pillars of quantum technologies which allow for developing a new generation of sensors for detecting gravitational, magnetic and electric fields as well as estimating physical quantities with unprecedented precision, well beyond the capability of conventional classical sensors. The application prospect of quantum sensors is very wide and includes mining Crawford et al. (2021), environmental monitoring  Lanzagorta et al. (2015), biological imaging Wu et al. (2016a); Taylor and Bowen (2016); Aslam et al. (2023), space exploration Kaltenbaek et al. (2021), radar technology Maccone and Ren (2020); Karsa et al. (2024); Casariego et al. (2023), searching for fundamental particles Brady et al. (2022); Backes et al. (2021); Kang et al. (2024); Wang et al. (2022a); Jiang et al. (2021a) — the list goes on. There are several reasons that make quantum sensors superior over their classical counterparts. First, the even smaller sizes of quantum sensors, which can be fabricated at the atomic scales, allows for excellent spatial resolution Schirhagl et al. (2014); Zhang et al. (2021). Second, the natural delicacy of quantum superposition allows quantum sensors to detect very weak signals Hong et al. (2024); Jiang et al. (2021b). Third, a combination of quantum particles can benefit from a unique property of quantum physics, namely entanglement, which allows for quantum enhanced precision in comparison with a classical sensor using the same amount of resources Giovannetti et al. (2004, 2011, 2006).
@@
Quantum criticality has already been identified as a resource for sensing in both equilibrium and non-equilibrium many-body systems. Originally, quantum phase transition in the transverse Ising model has been explored for metrology purposes Zanardi and Paunković (2006). It has been then extended to various types of criticalities including, second-order Zanardi and Paunković (2006); Zanardi et al. (2007); Gu et al. (2008); Zanardi et al. (2008); Invernizzi et al. (2008); Gu (2010); Gammelmark and Mølmer (2011); Skotiniotis et al. (2015); Rams et al. (2018); Wei (2019); Chu et al. (2021); Liu et al. (2021); Montenegro et al. (2021); Mirkhalaf et al. (2021); Di Candia et al. (2023); Mihailescu et al. (2024a), first-order Raghunandan et al. (2018); Heugel et al. (2019); Yang and Jacob (2019), topological Budich and Bergholtz (2020); Sarkar et al. (2022); Koch and Budich (2022); Yu et al. (2024a); Sarkar et al. (2024), non-Hermitian systems Chen et al. (2017); Yu et al. (2020); McDonald and Clerk (2020); Wiersig (2014); Hodaei et al. (2017); Liu et al. (2016); Langbein (2018); Chen et al. (2019); Wang et al. (2020); Zhang et al. (2019); Lau and Clerk (2018); Kononchuk et al. (2022); Wiersig (2020); Xiao et al. (2024); Yu et al. (2024b), and Stark phase transitions He et al. (2023); Yousefjani et al. (2023a, 2024a). Apart from these distinct types of criticalities at equilibrium, non-equilibrium quantum phase transitions can also demonstrate quantum enhanced sensitivity. This includes dissipative Fernández-Lorenzo and Porras (2017); Baumann et al. (2010); Baden et al. (2014); Klinder et al. (2015); Rodriguez et al. (2017); Fitzpatrick et al. (2017); Fink et al. (2017); Ilias et al. (2022, 2024); Alipour et al. (2014), Floquet Mishra and Bayat (2021, 2022) and time crystal Montenegro et al. (2023); Cabot et al. (2024); Lyu et al. (2020); Iemini et al. (2024a); Yousefjani et al. (2024b) phase transitions. Each type of quantum criticality comes with its own features and characteristics. For instance, second-order quantum phase transitions take place in the ground state of a many-body system and are accompanied by spontaneous symmetry breaking and long-range correlations which are described by Landau-Ginzburg formalism Sachdev (1999). On the other hand, topological phase transitions neither show symmetry breaking nor follow the Landau-Ginzburg theorem Xu (2012) and Stark localization takes place across the entire spectrum and not just the ground state van Nieuwenburg et al. (2019); Schulz et al. (2019). While different types of quantum criticalities have their own characteristics, one feature is common among all of them for showing quantum enhanced sensitivity which is energy/quasi-energy gap closing. This is an interesting observation which guides us in searching for finding new potential quantum sensors with the capability of achieving quantum enhanced precision. Interestingly, the realization of criticality-based quantum sensors are becoming viable in various physical platforms, including nuclear magnetic resonance systems Liu et al. (2021) and Rydberg atoms Ding et al. (2022).
@@
Quantum criticality is in general a property of equilibrium states such as the Gibbs thermal state (which becomes the ground state at zero temperature limit) Sachdev and Safari (2011); Vojta (2003). Preparing such equilibrium states can be challenging in practice, demanding ultra-low temperatures or extremely long preparation times. In addition, criticality-enhanced quantum precision is only achievable around the phase transition point. This requires fine tuning and thus is mostly beneficial in the context of local sensing, where significant prior information about the parameter of interest is available. Beyond criticalities, the sensing capability of many-body systems have also been investigated in their strongly correlated phases, even in the presence of thermal fluctuations Mihailescu et al. (2023, 2024b, 2024a). Moreover, as an alternative approach, one can also use non-equilibrium dynamics of many-body systems for sensing unknown parameters. These setups are more experimental friendly as their initial state can be a simple product state and entanglement is generated naturally during the dynamics. Non-equilibrium dynamics in many-body systems can be induced through different methods, including quantum quench Mitra (2018); Daley et al. (2012); Calabrese and Cardy (2006); Calabrese et al. (2011); Shi et al. (2024); Gietka et al. (2022), measurement quench Pouyandeh et al. (2014); Bayat et al. (2018); Skinner et al. (2019a); Li et al. (2018a), external driving Chu and Telnov (2004); Gritsev and Polkovnikov (2017); Oka and Kitamura (2019) and adiabatic evolution Mihailescu et al. (2024b). Apart from the probe size, in non-equilibrium quantum sensors, time is also a key resource which plays an important role and has to be included in the definition of standard limit and quantum enhanced sensitivity. In general, Fisher information scales as $I\sim t^{\alpha}L^{\beta}$, where $t$ is time, $L$ is the system size and $\alpha$ and $\beta$ are two exponents Ilias et al. (2022). In non-equilibrium sensors, the standard limit is defined as $\alpha{=}\beta{=}1$. Any situation with $\alpha>1$ or $\beta>1$ is called quantum enhanced sensitivity with the special case of Heisenberg limit, defined as $\alpha{=}\beta{=}2$. In certain systems in which the unitary dynamics preserves the number of excitations, the Fisher information may scale with the number of excitations too. In fact, in Ref. Manshouri et al. (2024) it was shown that in excitation-preserving dynamics Fisher information scales as $I\sim t^{\alpha}L^{\beta}N^{\gamma}$, where $N$ is the number of excitations and $\gamma$ is its corresponding exponent. This is an interesting observation which shows how the sensitivity increases from single excitation subspace, i.e. $N{=}1$, to half filling, i.e. $N{=}L/2$. In addition, Ref. Gietka and Ritsch (2023) proposes using spin-orbit coupled quantum gases to overcome the Heisenberg limit and Ref. Gietka et al. (2022) demonstrates that Fisher information can exhibit exponential scaling as a physical effect by quenching dynamics far beyond the critical point. Unlike criticality-based quantum sensors, the criteria for achieving quantum enhanced precision is not yet well-characterized in non-equilibrium quantum many-body probes. For instance, we know that localized phases are not good for sensing He et al. (2023); Yousefjani et al. (2023a, 2024a); Sahoo et al. (2024) but what features in extended phases are crucial for quantum enhanced sensitivity are not yet identified.
@@
In practice, no quantum device is perfectly isolated from its environment and thus decoherence is an inevitable part of any quantum protocol. Consequently, any quantum dynamics is actually an open system evolution. Different formalisms have been developed for addressing open system dynamics, among them are: (i) master equation Carmichael (1993); and (ii) non-Hermitian Hamiltonians Rotter (2009). In the former, the evolution of the system and its environment is described by a unitary operation. By tracing out the environment, the evolution of the system, typically within Born-Markov approximations, is thus characterized by the Lindblad (or Gorini-Kossakowski-Sudarshan-Lindblad) master equation Breuer et al. (2002). In this formalism a quantum system initially prepared in a pure state may indeed evolve to a mixed state due to the action of jump operators — a signature of information leakage (irreversible process) to the environment. In non-Hermitian formalism, the environment is monitored for tracking a trajectory in which no jump operator action takes place Plenio and Knight (1998); Daley (2014). As a result, an initially pure quantum state remains pure throughout the dynamics, despite being evolved effectively as an open quantum system. Other mechanisms that address or profit from dissipation include decoherence-free subspaces Lidar et al. (1998); Duan and Guo (1997), by coupling the quantum resources to a common environment that can be measured at least in part Braun and Martin (2011) and coherent averaging (Fraïsse and Braun, 2015) more generally leading to quantum-enhanced measurements without entanglement (Braun et al., 2018), dynamical decoupling Viola et al. (1999); Khodjasteh and Lidar (2005), and quantum reservoir engineering Poyatos et al. (1996). The latter technique involves adding an engineered Lindbladian term to the open system dynamics. This approach can be designed to steer the system into a desired pure steady-state, effectively cooling the system to that state Myatt et al. (2000); Pielawa et al. (2007, 2010); Groszkowski et al. (2022). Both master equation and non-Hermitian formalisms have been investigated for quantum sensing. Indeed, a general joint (system plus environment) quantum Fisher information upper bound has been formulated Escher et al. (2011a). Hence, one would expect that such open quantum system dynamics may also achieve quantum enhanced sensitivity provided that the system goes through a phase transition, which is again accompanied by an eigenvalue gap closing. In the master equation formalism, the phase transition is determined for the steady state of the system at the point that the eigenvalues of the Liouvillian operator shows gap closing Minganti et al. (2018). An example of such systems with quantum enhanced sensitivity is boundary time crystals Montenegro et al. (2023). In the non-Hermitian systems, quantum enhanced sensitivity has been shown to emerge at the exceptional points Wiersig (2014); Kononchuk et al. (2022), where two or more eigenvalues and corresponding eigenstates coalesce, as well as in the cases of different types of gap closing in the complex energy plane Sarkar et al. (2024).
@@ II.1 Global and local estimation theory
A suitable strategy to estimate a parameter usually depends on the amount of a priori information one has on the parameter itself. If no a priori information is avalaible, the estimation strategy, i.e. the choice of the measure to be performed on the probe and the procedure employed to process data, should be optimal in average i.e. should be equally good for any value of the parameter. The set of techniques and methods used to choose and optimize the figure of merit in this case is known as global estimation theory Helstrom (1976); Holevo (2011a). Once a global estimation has been performed an experimenter is left with some a priori information to be improved, i.e. one should look for figures of merits that select the best estimation strategy for a specific value or range of values of a parameter. This optimization problem is addressed by local estimation theory Bengtsson and Życzkowski (2017); Van Trees (2004); Helstrom (1976) which allows one to achieve the ultimate bound on precision. The quantum version, namely the local quantum estimation theory Braunstein and Caves (1994); Paris (2009) is the fundamental tool of quantum probing, sensing and metrology and is the main subject of the present review, with the exception of Section XIV, where various aspects of global sensing are discussed.
@@ II.3 Optimizing the classical bound: the QFI
Note that the derivative of the quantum state satisfies other expressions, e.g., the Right Logarithmic Derivative (RLD) Albarelli et al. (2020). This gives rise to different quantum lower bounds. Hence, as opposed to the CFI, the QFI is not unique (see Ref. Meyer (2021) for a thorough discussion and Ref. Albarelli et al. (2020) for different achievable bounds).
The QFI of Eq. (6) accounts for the ultimate precision limit for the estimation of $\theta$. Note that the QFI, as the maximization of the CFI over all possible measurements, depends on the quantum statistical model $\rho_{\theta}$ and the SLD $L_{\theta}$ only, and it does not depend explicitly on a particular measurement basis. The optimal POVM can be composed with the set of projectors over the eigenstates of $L_{\theta}$ with an associated optimal quantum estimatorParis (2009).
To saturate the quantum Cramér-Rao bound in Eq. (2), one needs both the optimal POVM (as determined by the support of $L_{\theta}$) and an optimal classical estimator for post-processing data, such as the maximum-likelihood estimator in the limit of large data sets Olivares and Paris (2009); Pezzé et al. (2007); Hradil et al. (1996); Le Cam (1986); Cramer (1946); Braunstein and Caves (1994); Helstrom (1967) (refer to Refs. Rubio and Dunningham (2019, 2020) for Bayesian analysis in the scenario of limited data). The optimal POVM ensures that the CFI equals the QFI, i.e., $I(\theta)=I_{Q}(\theta)$. This is crucial as it determines that the single-parameter quantum Cramér-Rao bound can always be achieved, provided that both the estimator and the measurement basis are optimal.
@@
The QFI of Eqs. (6)-(11) applies to any quantum state. However, for pure quantum states $\rho_{\theta}=|\psi_{\theta}\rangle\langle\psi_{\theta}|$ such that $\text{Tr}[\rho_{\theta}^{2}]=1$, SLD $L_{\theta}=2\partial_{\theta}\rho_{\theta}$ and the QFI adopts a simpler form given by Paris (2009); Sidhu and Kok (2020); Liu et al. (2019):